OpenStudy (anonymous):

So, I need some help with differential equations. I have \[y_1'=-6y_1+7y_2\]\[y_2'=-6y_1+15y_2\], and with the conditions \[\left(\begin{matrix}y_1(0) \\ y_2(0)\end{matrix}\right)=\left(\begin{matrix}5 \\ 0\end{matrix}\right)\] Then I need to find the 2nd coordinatfunction. \[y_2(t)=2e^{-3t}-Oe^{\lambda_2t}\] This means that I need to find the value for O.

OpenStudy (anonymous):

I have in previous questions, found the eigenvectors and values for this system. \[\lambda_1=-3 \]With eigenvector:\[u=\left(\begin{matrix}1 \\ 2\end{matrix}\right) \]\[\lambda_2=12\] With the eigenvector\[v=\left(\begin{matrix}3 \\ 1\end{matrix}\right)\]

OpenStudy (anonymous):

In case question's math doesnt load: (it doesnt for me): So, I need some help with differential equations. I have \[y_1'=-6y_1+7y_2\]\[y_2'=-6y_1+15y_2\], and with the conditions \[\left(\begin{matrix}y_1(0) \\ y_2(0)\end{matrix}\right)=\left(\begin{matrix}5 \\ 0\end{matrix}\right)\] Then I need to find the 2nd coordinatfunction. \[y_2(t)=2e^{-3t}-Oe^{\lambda_2t}\] This means that I need to find the value for O.

OpenStudy (michele_laino):

here we can rewrite such system of ODEs, like below: \[\left( {\begin{array}{*{20}{c}} {{y_1}'} \\ {{y_2}'} \end{array}} \right) = \left( {\begin{array}{*{20}{c}} { - 6}&7 \\ { - 6}&{15} \end{array}} \right) \cdot \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = A\left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right)\] so, we have to find the eigenvalues of A, first

OpenStudy (anonymous):

Is it only me having trouble loading the site sometimes?

OpenStudy (michele_laino):

please wait a moment, I compute the eigenvalues of A

OpenStudy (anonymous):

ohh, wait. It is supposed to be 9, not 7. In \[y_1'=-6y_1+9y_2\]

OpenStudy (anonymous):

Not that it matters that much, just need to understand the general way to do it

OpenStudy (michele_laino):

ok! Then the eigenvalues are given by the subsequent equation: \[\Large \begin{gathered} \det \left( {\begin{array}{*{20}{c}} { - 6 - \lambda }&9 \\ { - 6}&{15 - \lambda } \end{array}} \right) = 0 \Rightarrow {\lambda ^2} - 9\lambda + 36 = 0 \hfill \\ \hfill \\ {\lambda _1} = - 3,\quad {\lambda _2} = 12 \hfill \\ \end{gathered} \] and the solution is: \[\Large \begin{gathered} {y_1} = {c_1}{e^{ - 3t}} + {c_2}{e^{12t}} \hfill \\ {y_2} = {c_3}{e^{ - 3t}} + {c_4}{e^{12t}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

oops.. I have made a typo: \[\Large \det \left( {\begin{array}{*{20}{c}} { - 6 - \lambda }&9 \\ { - 6}&{15 - \lambda } \end{array}} \right) = 0 \Rightarrow {\lambda ^2} - 9\lambda - 36 = 0\]

OpenStudy (anonymous):

Yes, I have found out so much. So now I need to find the values for \[c_4\]

OpenStudy (michele_laino):

now we have to apply the initial condition, so we can write this: \[\Large \left\{ \begin{gathered} 5 = {c_1} + {c_2} \hfill \\ 0 = {c_3} + {c_4} \hfill \\ \end{gathered} \right.\]

OpenStudy (anonymous):

2 equations with 4 variables?

OpenStudy (michele_laino):

please sorry I have made an error, here are the right equations:

OpenStudy (anonymous):

Well, actually I got it now. Because I have given \[c_3\] That would leave \[0=2-c_4<->c_4=2\]

OpenStudy (anonymous):

But if I didnt have \[c_3\] would I still be able to solve all the 4 variables with 2 equations?

OpenStudy (phi):

there are only 2 unknown constants

OpenStudy (anonymous):

Are they the same, \[c_1=c_3\] and \[c_2=c_4\]?

OpenStudy (phi):

also, I got eigenvalue 12 going with vector <1 2> and -3 with <3 1>

OpenStudy (anonymous):

Thats the same as I did @phi

OpenStudy (michele_laino):

here are the solutions: \[\Large \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 1 \\ 2 \end{array}} \right){e^{ - 3t}} + {c_2}\left( {\begin{array}{*{20}{c}} 3 \\ {21} \end{array}} \right){e^{12t}}\]

OpenStudy (phi):

you can think of the answer as \[ y= c_1 x_1 \exp(\lambda_1 t) + c_2 x_2 \exp(\lambda_2 t)\] where y, x1 and x2 are all vectors

OpenStudy (anonymous):

ah, yea.

OpenStudy (anonymous):

Could you help me with a follow up question to this?

OpenStudy (michele_laino):

sorry another typo: \[\Large \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 1 \\ 2 \end{array}} \right){e^{ - 3t}} + {c_2}\left( {\begin{array}{*{20}{c}} 3 \\ 1 \end{array}} \right){e^{12t}}\]

OpenStudy (anonymous):

Now I have:\[y_1(0)=1\] and\[\lim_{t \rightarrow \infty}y_1(t)=0\] I need to find the missing values: \[y_2(t)=\frac{ 1 }{ 3 }e^{-3t}-(missing)e^{12t}\]

OpenStudy (phi):

double check. I think it's \[ \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 1 \\ 2 \end{array}} \right){e^{ 12t}} + {c_2}\left( {\begin{array}{*{20}{c}} 3 \\ 1 \end{array}} \right){e^{-3t}}\]

OpenStudy (anonymous):

Yes, that is correct @phi when I insert the eigenvalue -3 I get the vector (3,1)

OpenStudy (michele_laino):

yes! I got the same result. So the solution, is: \[\Large \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 3 \\ 1 \end{array}} \right){e^{ - 3t}} + {c_2}\left( {\begin{array}{*{20}{c}} 1 \\ 2 \end{array}} \right){e^{12t}}\] now, we have to set \(c_2=0\), if we want \[{y_1} \to 0\]

OpenStudy (michele_laino):

when t---> +infinity

OpenStudy (anonymous):

I am not sure I follow, \[c_2\] in which equation?

OpenStudy (michele_laino):

sorry, I meant in my last equation

OpenStudy (michele_laino):

\[\Large \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 3 \\ 1 \end{array}} \right){e^{ - 3t}} + {c_2}\left( {\begin{array}{*{20}{c}} 1 \\ 2 \end{array}} \right){e^{12t}}\]

OpenStudy (michele_laino):

since we have: \[\mathop {\lim }\limits_{x \to + \infty } {e^{12t}} = + \infty \]

OpenStudy (anonymous):

We have \[\lim_{t \rightarrow \infty}y_1(t)=0\]

OpenStudy (michele_laino):

since \[\Large {e^{ - 3t}},\;\;{e^{12t}}\] are two linearly independent functions, we have to set \(c_2=0\), otherwise, we wiil not get such condition you wrote

OpenStudy (michele_laino):

will*

OpenStudy (anonymous):

Yes, how do I translate that into \[y_2(t)\]?

OpenStudy (michele_laino):

we get this: \[\Large \left( {\begin{array}{*{20}{c}} {{y_1}} \\ {{y_2}} \end{array}} \right) = {c_1}\left( {\begin{array}{*{20}{c}} 3 \\ 1 \end{array}} \right){e^{ - 3t}} = \left( {\begin{array}{*{20}{c}} {3{c_1}{e^{ - 3t}}} \\ {{c_1}{e^{ - 3t}}} \end{array}} \right)\] now, I can write: \[\Large 3{c_1} = 1 \Rightarrow {c_1} = \frac{1}{3}\]

OpenStudy (michele_laino):

and, of course: \[\Large {\text{missing value}} = {c_2} = 0\]

OpenStudy (anonymous):

ah, yea. Thank you SO much :)

OpenStudy (michele_laino):

:)